# Prosoluble subgroups of the profinite completion of the fundamental group of compact 33 33-manifolds

### v2

@@ 2.2 The fundamental group of a finite graph of pro-$\mathcal{C}$ groups
For a totally new construction of the fundamental group of a profinite graph of pro-$\mathcal{C}$ groups using an approach similar to the abstract case, the reader should check [AZ22].
@@
For the general case, given any open normal subgroup $U$ of $G$, we have (see [AZ22, Proposition 2.15 and Theorem 3.9])
@@
With standard arguments and can deduce the following:
@@
Given a finite graph of pro-$\mathcal{C}$ groups $(\mathcal{G},\Gamma)$ with fundamental group $G$, the correspondent standard graph is
@@ 2.3 Auxiliary Results
Put $T=H\Sigma$. Since the quotient of $T$ modulo every open subgroup $U$ of $H$ is connected, $T$ is connected. By [Rib17, Proposition 3.7.3], $\Sigma$ is simply connected. By [Rib17, Proposition 3.9.2], $\Delta=H\backslash T$ is simply connected and so is a finite tree. By [Rib17, Theorem 6.6.1], $H=\Pi_{1}(\mathcal{H},\Delta)$ whose edge and vertex groups are stabilizers of the corresponding edges and vertices of a connected transversal of $\Delta$ in $T$. ∎
@@
The previous result does not holds if there is some $e$ such that $|G_{v}:G_{e}|=|G_{w}:G_{e}|=2$. When it occurs we will say that $\left\langle G_{v},G_{e},G_{w}\right\rangle$ is of dihedral type or that the edge $e$ generates a dihedral type group.
@@
Let $\mathcal{C}$ be a variety does not containing all finite groups and $G$ a pro-$\mathcal{C}$ group acting on a simply connected profinite graph $\Gamma$. Suppose that $G$ has no edges $e$ generating a dihedral type group. Then for every edge $e$ with vertices $v,w$ and non-trivial edge stabilizer, we have $G_{v}=G_{e}$ or $G_{w}=G_{e}$.
@@
Suppose that, for a non-trivial edge stabilizer $e$ with vertices $v$ and $w$, we have $G_{v}$ and $G_{w}$ both different from $G_{e}$. Note that the index of $G_{e}$ in $G_{v}$ or in $G_{w}$ is greater than 22 22. By Proposition 2.15, $H=\left\langle G_{v},G_{w}\right\rangle$ is a not a pro-$\mathcal{C}$ subgroup of $G$, a contradiction. ∎
@@
Let $G=G_{1}\amalg_{H}G_{2}$ be a splitting of a profinite group $G$ as an amalgamated free profinite product of profinite groups $G_{1},G_{2}$ and $H_{1}\leq G_{1},H_{2}\leq G_{2}$ be subgroups such that $H_{1}\cap H\leq U\geq H\cap H_{2}$ for some open normal subgroup $U$ of $G$. Then $\langle H_{1},H_{2}\rangle=L_{1}\amalg_{K}L_{2}$ with $L_{1}\leq H_{1}U,L_{2}\leq H_{2}U,K\leq HU$.
@@
By [ZM89, Proposition 4.4], $\langle H_{1},H_{2}\rangle=L_{1}\amalg_{K}L_{2}$ with $L_{1}\leq G_{1}$, $L_{2}\leq G_{2}$, $K\leq H$. Recall that the fundamental group $\Pi_{1}(\mathcal{G},\Gamma)$ of a finite graph of profinite groups is a completion of the abstract fundamental group ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$ of the same graph of groups with respect to to the topology generated by the finite index normal subgroups $N$ of ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$ such that $N\cap\mathcal{G}(v)$ is open in $\mathcal{G}(v)$. Note that $G$ is the fundamental group of a finite graph of profinite groups. We can consider $U\cap G_{1}$ and $U\cap G_{2}$ to apply [CZ23, Lemma 3.1] on ${G}^{\text{abs}}=G_{1}\ast_{H}G_{2}$ so we have ${\left\langle H_{1},H_{2}\right\rangle}^{\text{abs}}=L_{1}\ast_{K}L_{2}$ where $L_{1}\leq H_{1}(U\cap G_{1})$, $L_{2}\leq H_{2}(U\cap G_{2})$ and $K\leq H\cap U$. Then $\left\langle H_{1},H_{2}\right\rangle=\overline{L}_{1}\amalg_{K}\overline{L}_{2}$ with $\overline{L}_{1}\leq H_{1}U$, $\overline{L}_{2}\leq H_{2}U$ and $K\leq HU$. ∎
@@
Let $G=G_{1}\amalg_{H}G_{2}$ be a splitting of the profinite group $G$ as an amalgamated free profinite product of profinite groups $G_{1},G_{2}$ and $H_{1}\leq G_{1},H_{2}\leq G_{2}$ be subgroups such that $H_{1}\cap H=1=H_{2}\cap H$. Then $\langle H_{1},H_{2}\rangle=H_{1}\amalg H_{2}$.
@@
Since $U$ in the preceding lemma is arbitrary, the result follows. ∎
@@ 2.4 Relatively projective groups
As was already mentioned in the introduction, the profinite analogue of the Kurosh Subgroup Theorem does not hold for free products of profinite groups and so the algebraic structure of closed subgroups of free profinite products is not clear. To compensate this, Dan Haran (see [Har87]) introduced the notion of relatively projective profinite groups and proved that a subgroup $H$ of a free profinite product $\coprod_{x\in X}G_{x}$ is projective relatively to the family $\{H\cap G_{x}^{g}\mid x\in X,g\in G\}$. We shall use a slightly more restricted notion here from [HZ23] as it better fits into the profinite version of Bass-Serre theory and suffices for our purpose.
@@
A group pile (or just pile) is a pair $\mathbf{G}=(G,T)$ consisting of a profinite group $G$, a profinite space $T$ and a continuous action of $G$ on $T$.
@@
A morphism of groups piles $\alpha:(G,T)\to(G^{\prime},T^{\prime})$ consists of a group homomorphism $\alpha:G\to G^{\prime}$ and a continuous map $\alpha:T\to T^{\prime}$ such that $\alpha(t^{g})=\alpha(t)^{\alpha(g)}$ for all $t\in T$ and $g\in G$. The above morphism is an epimorphism if $\alpha(G)=G^{\prime}$, $\alpha(T)=T^{\prime}$ and for every $t^{\prime}\in T^{\prime}$ there is $t\in T$ such that $\alpha(t)=t^{\prime}$ and $\alpha(G_{t})={G^{\prime}}_{t^{\prime}}$. It is rigid if $\alpha$ maps $G_{t}$ isomorphically onto ${G^{\prime}}_{\alpha(t)}$, for every $t\in T$, and the induced map of the orbit spaces $G\backslash T\to G^{\prime}\backslash T^{\prime}$ is a homeomorphism.
@@
An embedding problem for a pile $\mathbf{G}$ is a pair $(\varphi,\alpha)$ where $\varphi:\mathbf{G}\to\mathbf{H}$ and $\alpha:\mathbf{G}^{\prime}\to\mathbf{H}$ are morphism of group piles such that $\alpha$ is an epimorphism. It is finite if $\mathbf{G}^{\prime}$ is finite. It is rigid if $\alpha$ is rigid. A solution to the embedding problem is a morphism $\gamma:\mathbf{G}\to\mathbf{G}^{\prime}$ such that $\alpha\circ\gamma=\varphi$.
@@
A pile $\mathbf{G}$ is $\mathcal{C}$-projective if every finite rigid $\mathcal{C}$-embedding problem for $\mathbf{G}$ has a solution. A group $G$ is $\mathcal{G}$-projective if every finite $\mathcal{C}$-embedding problem has a weak solution.
It follows from [Zal23b, Remark 2.4] that if $\mathbf{G}$ is projective then $G$ is $\mathcal{G}$-projective.
@@
    Let $\mathcal{C}$ be a class of finite groups closed for subgroups, quotients and extensions. Let $(G,T)$ be a projective pro-$\mathcal{C}$ pile. Suppose that exists a continuous section $\sigma:G\backslash T\to T$. Then $G$ acts on a pro-$\mathcal{C}$ tree with trivial edge stabilizers and non-trivial vertex stabilizers being stabilizers of points in $T$.
We also need of the following proposition:
@@
[Zal23b, Proposition 2.5] Let $\mathcal{C}$ be a class of finite groups closed for subgroups. quotients and extensions. Let $(G,T)$ be a $\mathcal{C}$-projective pile such that there exists a continuous section $\sigma:G\backslash T\to T$. Then $G$ embeds into a free pro-$\mathcal{C}$ product
@@ 3 Acylindrical actions and prosoluble groups
In this section we prove a generalization of [WZ17, Lemma 11.2]. This result allows us to deduce that if $G$ acts $k$-acylindrically on a profinite simply connected graph $T$, then for any closed normal subgroups $N$ generated by vertex stabilizers the quotient group $G/N$ acts $k$-acylindrically on $N\backslash T$.
@@
We say that a finite graph of profinite groups $(\mathcal{G},\Gamma)$ is $k$-acylindrical if its fundamental group acts $k$-acylindrically on its standard pro-$\mathcal{C}$ tree.
@@
and $G_{e}\neq 1$. Put $v=d_{0}(e)$, $w=d_{1}(e)$ and set $H=\left\langle G_{v},G_{w}\right\rangle$. Let $D=H(v\cup e\cup w)$. Since for every open subgroup $U$ of $H$ the quotient of $U\backslash D$ is connected, so is $D$. By Lemma 2.14 $H=G_{v}\amalg_{G_{e}}G_{w}$. Since $G_{e}$ is normal in $H$, by [Rib17, Proposition 4.2.2, (b)], $G_{e}$ acts trivially on $D$, contradicting the $k$-acylindricity of the action. This finishes the proof. ∎
@@
Suppose $\mathcal{C}$ does not contain all finite groups and let $G$ be a pro-$\mathcal{C}$ group acting $k$-acylindrically on a simply connected profinite graph $\Gamma$. Then the maximal abstract subgraph $D$ of $\Gamma$ such that each $e\in D$ satisfies $G_{e}\neq 1$ has finite diameter $\leq 2k$. Moreover, for every connected component $D^{*}$ of $D$ there at most one other connected component at finite distance to $D^{*}$ in which case stabilizers of edges and vertices of $D^{*}$ are of order 22 22.
@@
Suppose on the contrary there exists a finite connected subgraph $\Sigma$ of $D$ whose diameter is $>2k$. Let $H=\left\langle G_{v}\mid v\in V(\Sigma)\right\rangle$ and $T=H\Sigma$. By Lemma 2.14 and [Rib17, Proposition 4.1.1], $H$ is the fundamental group of a tree of pro-$\mathcal{C}$ groups $\Pi_{1}(\mathcal{H},H\backslash T)$ whose edge and vertex groups are stabilizers of the corresponding edges and vertices of a connected transversal of $H\backslash T$ in $T$. Hence, by Proposition 3.4 there is no edge $e$ such that
@@
with $G_{e}\neq 1$. Let $(\mathcal{H}_{red},H\backslash T_{red})$ be a reduced tree of groups obtained from $(\mathcal{H},H\backslash T)$. If it has more than one vertex, then $H$ cannot be pro-$\mathcal{C}$ by Proposition 2.12. Hence, $(\mathcal{H}_{red},H\backslash T_{red})$ has one vertex $v$ only which implies that $H$ fixes a vertex $w$ in $T$ and, in fact, $w\in\Sigma$. As the diameter of $\Sigma$ is greater than $2k$, it follows that there exists a geodesic $[u,w]$ in $\Sigma$ of length greater than $k$ which is fixed by $H_{u}\neq 1$ (see [Rib17, Corollary 4.1.6]) contradicting the $k$-acylindricity of the action. This proves the first part of the statement.
@@
Suppose whithin the hypothesis of Theorem 3.5 $\Gamma$ has infinite diameter. Then $\Gamma$ possesses an edge $e$ with $G_{e}=1$, namely $G_{e}=1$ for any $e\not\in D$.
@@
Suppose that $G\backslash D$ has finitely many connected components. Then $G$ acts on a simply connected profinite $G$-quotient graph $T$ with trivial edge stabilizers and vertex stabilizers equal to the stabilizers of some vertices $\Gamma$. In particular, $G$ has only a finite number of maximal stabilizers of vertices in $T$ up to conjugation and this number equals to the number maximal vertex stabilizers in $\Gamma$ that does not exceed the number of connected components of $G\backslash D$.
@@
By [CZ22, Lemma 3.2] the closure of any connected component $\overline{D^{*}}$ is a profinite subgraph of $\Gamma$ of diameter $\leq 2k$, hence $\overline{D}=\bigcup_{i}D_{i}^{*}$ has finite diameter, where $D_{i}^{*}$ runs via orbit representatives of the connected components of $D$. Moreover, as every connected component $\overline{D}_{\ast}$ of $\overline{D}$ has finite diameter, its stabilizer $\operatorname{Stab}(\overline{D}_{\ast})$ fixes some vertex $v_{\ast}\in V(\overline{D}_{\ast})$. Then collapsing all connected components of $\overline{D}$ we obtain a $G$-invariant simply connected quotient graph $T$ of $\Gamma$ (see [Rib17, Proposition 3.9.1]) on which $G$ acts with trivial edge stabilizers and only non-trivial vertex stabilizers of $T$ are stabilizers of connected component $\operatorname{Stab}(\overline{D}_{\ast})=G_{v_{\ast}}$ of $\overline{D}$. This concludes the proof. ∎
Let $G$ be the fundamental group of a finite graph of profinite groups and $S$ its standard graph. For any normal open subgroup $U$ of $G$ we set
$\tilde{U}=\left\langle U\cap H_{v}:v\in V(S)\right\rangle.$

### v3

@@ 2.2 The fundamental group of a finite graph of pro-$\mathcal{C}$ groups
For a totally new construction of the fundamental group of a profinite graph of pro-$\mathcal{C}$ groups using an approach similar to the abstract case, the reader should check [AZ22]. Here we will use the standard construction done in [Rib17].
@@
For the general case, given any open normal subgroup $U$ of $G$, consider
We have (see [AZ22, Proposition 2.15 and Theorem 3.9])
@@
With standard arguments one can deduce the following:
@@
Given a finite graph of pro-$\mathcal{C}$ groups $(\mathcal{G},\Gamma)$ with fundamental group $G$, the corresponding standard graph is
@@ 2.3 Auxiliary Results
Put $T=H\Sigma$. Since the quotient of $T$ modulo every open subgroup $U$ of $H$ is connected, $T$ is connected. By [Rib17, Proposition 3.7.3], $\Sigma$ is simply connected. By [Rib17, Proposition 3.9.2], $\T\displaystyle\Delta=H\backslash$ T is simply connected and so is a finite tree. By [Rib17, Theorem 6.6.1], $H=\Pi_{1}(\mathcal{H},\Delta)$ whose edge and vertex groups are stabilizers of the corresponding edges and vertices of a connected transversal of $\Delta$ in $T$. ∎
@@
The previous result does not hold if there is some $e$ such that $|G_{v}:G_{e}|=|G_{w}:G_{e}|=2$. When it occurs we will say that $\left\langle G_{v},G_{e},G_{w}\right\rangle$ is of dihedral type or that the edge $e$ generates a dihedral type group.
@@
Let $\mathcal{C}$ be a variety not containing all finite groups and $G$ a pro-$\mathcal{C}$ group acting on a simply connected profinite graph $\Gamma$. Suppose that $G$ has no edges $e$ generating a dihedral type group. Then for every edge $e$ with vertices $v,w$ and non-trivial edge stabilizer, we have $G_{v}=G_{e}$ or $G_{w}=G_{e}$.
@@
Suppose that, for a non-trivial edge stabilizer $e$ with vertices $v$ and $w$, we have $G_{v}$ and $G_{w}$ both different from $G_{e}$. By hypothesis, the index of $G_{e}$ in $G_{v}$ or in $G_{w}$ is greater than $2$. By Proposition 2.15, $H=\left\langle G_{v},G_{w}\right\rangle$ is a not a pro-$\mathcal{C}$ subgroup of $G$, a contradiction. ∎
@@
Let $G=G_{1}\amalg_{H}G_{2}$ be a splitting of a profinite group $G$ as an amalgamated free profinite product of profinite groups $G_{1},G_{2}$ and $H_{1}\leq G_{1},H_{2}\leq G_{2}$ be subgroups such that $H_{1}\cap H\leq U\geq H\cap H_{2}$ for some open normal subgroup $U$ of $G$. Then $\langle H_{1},H_{2}\rangle=L_{1}\amalg_{K}L_{2}$ with $L_{1}\leq H_{1}U,L_{2}\leq H_{2}U,K\leq H\cap U$.
@@
By [ZM89a, Proposition 4.4], $\langle H_{1},H_{2}\rangle=L_{1}\amalg_{K}L_{2}$ with $L_{1}\leq G_{1}$, $L_{2}\leq G_{2}$, $K\leq H$. Recall that the fundamental group $\Pi_{1}(\mathcal{G},\Gamma)$ of a finite graph of profinite groups is a completion of the abstract fundamental group ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$ of the same graph of groups with respect to to the topology generated by the finite index normal subgroups $N$ of ${\Pi_{1}}^{\text{abs}}(\mathcal{G},\Gamma)$ such that $N\cap\mathcal{G}(v)$ is open in $\mathcal{G}(v)$. Note that $G$ is the fundamental group of a finite graph of profinite groups. We can consider $U\cap G_{1}$ and $U\cap G_{2}$ to apply [CZ23, Lemma 3.1] on ${G}^{\text{abs}}=G_{1}\ast_{H}G_{2}$ to deduce that ${\left\langle H_{1},H_{2}\right\rangle}^{\text{abs}}=L_{1}\ast_{K}L_{2}$ where $L_{1}\leq H_{1}(U\cap G_{1})$, $L_{2}\leq H_{2}(U\cap G_{2})$ and $K\leq H\cap U$. Then $\left\langle H_{1},H_{2}\right\rangle=\overline{L}_{1}\amalg_{K}\overline{L}_{2}$ with $\overline{L}_{1}\leq H_{1}U$, $\overline{L}_{2}\leq H_{2}U$ and $K\leq H\cap U$. ∎
@@
Let $G=G_{1}\amalg_{H}G_{2}$ be a splitting of a profinite group $G$ as an amalgamated free profinite product of profinite groups $G_{1},G_{2}$ and $H_{1}\leq G_{1},H_{2}\leq G_{2}$ be subgroups such that $H_{1}\cap H=1=H_{2}\cap H$. Then $\langle H_{1},H_{2}\rangle=H_{1}\amalg H_{2}$.
@@
Since we have $H_{1}\cap H=1=H_{2}\cap H$ we deduce from Lemma 2.17 that $K\leq U$, $L_{1}\leq H_{1}U$, $L_{2}\leq H_{2}U$ for an arbitrary open normal subgroup $U$. Since the intersection of all such $U$ is trivial, $L_{1}=H_{1}$, $L_{2}=H_{2}$, $K=1$ and the result follows. ∎
Let $G$ be a pro-$\mathcal{C}$ group acting on a pro-$\mathcal{C}$ tree $T$. Recall that, for any subgroup $U$ of $G$, we set
$\widetilde{U}=\left\langle U\cap G_{v}:v\in V(T)\right\rangle.$
Proposition 2.19.
Let $G$ be a pro-$\mathcal{C}$ group acting on a pro-$\mathcal{C}$ tree $T$ and $U$ an open normal subgroup of $G$. Then $D_{U}=\{m\in\widetilde{U}\backslash T:(G/\widetilde{U})_{m}\neq 1\}$ is a profinite subgraph of $\T\displaystyle\widetilde{U}\backslash$ T.
Proof.
Put $G_{U}=G/\widetilde{U}$, $\T\displaystyle T_{U}=\widetilde{U}\backslash$ T. It suffices to prove that $D_{U}$ is closed in $T_{U}$. As $\{(G_{U})_{m}:m\in T\}$ is a continuous family (see [Rib17, Lemma 5.2.2]), the set $\{m\in\tilde{U}\backslash T:(G/\widetilde{U})_{m}=1\}=\{m\in\tilde{U}\backslash T:(G/\widetilde{U})_{m}\leq U/\widetilde{U}\}$ is open. Hence $D_{U}=\{m\in\widetilde{U}\backslash T:(G/\widetilde{U})_{m}\neq 1\}$ is closed. ∎
@@ 2.4 Relatively projective groups
As was already mentioned in the introduction, the profinite analogue of the Kurosh Subgroup Theorem does not hold for free profinite products of profinite groups and so the algebraic structure of closed subgroups of free profinite products is not clear. To compensate this, Dan Haran (see [Har87]) introduced the notion of relatively projective profinite groups and proved that a subgroup $H$ of a free profinite product $\coprod_{x\in X}G_{x}$ is projective relatively to the family $\{H\cap G_{x}^{g}\mid x\in X,g\in G\}$. We shall use a slightly more restricted notion here from [HZ23] as it better fits into the profinite version of Bass-Serre theory and suffices for our purpose.
@@
A group-pile is a pair $\mathbf{G}=(G,T)$ consisting of a profinite group $G$, a profinite space $T$ and a continuous action of $G$ on $T$. We say that $\mathbf{G}$ is finite if both $G$ and $T$ are finite.
@@
[…]
@@
An embedding problem for a group-pile $\mathbf{G}$ is a pair $(\varphi,\alpha)$ where $\varphi:\mathbf{G}\to\mathbf{H}$ and $\alpha:\mathbf{G}^{\prime}\to\mathbf{H}$ are morphisms of group-piles such that $\alpha$ is an epimorphism. It is finite if $\mathbf{G}^{\prime}$ is finite. It is rigid if $\alpha$ is rigid. A solution to the embedding problem is a morphism $\gamma:\mathbf{G}\to\mathbf{G}^{\prime}$ such that $\alpha\circ\gamma=\varphi$.
@@
For the next definition, let $\mathcal{G}$ be a continuous family of subgroups of $G$ and $\mathcal{C}$ a class of finite groups closed for subgroups, quotients and extensions.
@@
A pile $\mathbf{G}$ is $\mathcal{C}$-projective if every finite rigid $\mathcal{C}$-embedding problem for $\mathbf{G}$ has a solution. A group $G$ is $\mathcal{G}$-projective if every finite $\mathcal{C}$-embedding problem has a weak solution, that is, there exists a group homomorphism $\gamma:G\to G^{\prime}$ such that $\alpha\circ\gamma=\varphi$.
@@
    Let $\mathcal{C}$ be a class of finite groups closed for subgroups, quotients and extensions. Let $(G,T)$ be a projective pro-$\mathcal{C}$ pile. Suppose that exists a continuous section $\TT\displaystyle\sigma:G\backslash$ $\to T$. Then $G$ acts on a pro-$\mathcal{C}$ tree with trivial edge stabilizers and non-trivial vertex stabilizers being stabilizers of points in $T$.
We also need the following proposition:
@@
[Zal23b, Proposition 2.5] Let $\mathcal{C}$ be a class of finite groups closed for subgroups, quotients and extensions. Let $(G,T)$ be a $\mathcal{C}$-projective group-pile such that there exists a continuous section $\TT\displaystyle\sigma:G\backslash$ $\to T$. Then $G$ embeds into a free $\mathcal{C} pro$duct
@@ 3 Acylindrical actions and prosoluble groups
In this section we prove a generalization of [WZ17, Lemma 11.2]. This result allows us to deduce that if $\mathcal{C}$ does not contain all finite groups and if $G$ is a pro-$\mathcal{C}$ group acting $k$-acylindrically on a simply connected profinite graph $T$, then for any closed normal subgroup $N$ generated by vertex stabilizers the quotient group $G/N$ acts $k$-acylindrically on $\T\displaystyle N\backslash$ T.
@@
We say that a finite graph of profinite groups $(\mathcal{G},\Gamma)$ is $k$-acylindrical if its profinite fundamental group acts $k$-acylindrically on its standard profinite tree.
@@
and $G_{e}\neq 1$. Put $v=d_{0}(e)$, $w=d_{1}(e)$ and set $H=\left\langle G_{v},G_{w}\right\rangle$. Let $D=H(v\cup e\cup w)$. Since for every open subgroup $U$ of $H$ the quotient of $\D\displaystyle U\backslash$ D is connected, so is $D$. By Lemma 2.14 $H=G_{v}\amalg_{G_{e}}G_{w}$. Since $G_{e}$ is normal in $H$, by [Rib17, Proposition 4.2.2, (b)], $G_{e}$ acts trivially on $D$, contradicting the $k$-acylindricity of the action. This finishes the proof. ∎
@@
Recall that a profinite graph $\Gamma$ is also an abstract graph. Let $G$ be a profinite group acting on it. Put $D=\{e\cup d_{0}(e)\cup d_{1}(e)\mid G_{e}\neq 1\}$. Then $D$ is an abstract subgraph of $\Gamma$ and is maximal among abstract subgraphs of $\Gamma$ having non-trivial edge stabilizers for all its edges.
@@
Suppose $\mathcal{C}$ does not contain all finite groups and let $G$ be a pro-$\mathcal{C}$ group acting $k$-acylindrically on a simply connected profinite graph $\Gamma$. Then the maximal abstract subgraph $D$ of $\Gamma$ such that each $e\in D$ satisfies $G_{e}\neq 1$ has finite diameter $\leq 2k$. Moreover, for every connected component $D^{*}$ of $D$ there at most one other connected component at a finite distance to $D^{*}$ in which case stabilizers of edges and vertices of $D^{*}$ are of order $2$.
@@
[…]
@@
[…]
@@
Suppose we are in the situation of Theorem 3.5 and $\Gamma$ has infinite diameter. Then $\Gamma$ possesses an edge $e$ with $G_{e}=1$, namely $G_{e}=1$ for any $e\not\in D$.
@@
Suppose that $\D\displaystyle G\backslash$ D has finitely many connected components. Then $G$ acts on a simply connected profinite $G$-quotient graph $T$ of $\Gamma$ with trivial edge stabilizers and vertex stabilizers equal to the stabilizers of some vertices $\Gamma$. Moreover, the number of maximal (by inclusion) $G$-stabilizers of vertices in $T$, up to conjugation, equals to the number of maximal $G$-stabilizers of vertices in $\Gamma$ and this number is bounded by the number of connected components of $\D\displaystyle G\backslash$ D.
@@
By [CZ22, Lemma 3.2] the closure of any connected component $\overline{D^{*}}$ is a profinite subgraph of $\Gamma$ of diameter $\leq 2k$, hence $\overline{D}=\bigcup_{i}D_{i}^{*}$ has finite diameter, where $D_{i}^{*}$ runs via orbit representatives of the connected components of $D$. Moreover, as every connected component $\overline{D}_{\ast}$ of $\overline{D}$ has finite diameter, its stabilizer $\operatorname{Stab}(\overline{D}_{\ast})$ fixes some vertex $v_{\ast}\in V(\overline{D}_{\ast})$ (cf. [Ser02, Corollary, page 20]). Now we can collapse all connected components of the profinite subgraph $\overline{D}$ of $\Gamma$ (see [Rib17, Exercise 2.1.11]; the resulting profinite quotient graph $T$ of $\Gamma$ is $G$-invariant and by [Rib17, Proposition 3.9.1] is simply connected. The edge stabilizers of $T$ are trivial; the vertices of $T$ whose stabilizers are non-trivial are exactly the vertices to which the connected components of $D$ were collapsed; thus the only non-trivial vertex stabilizers of $T$ are stabilizers of the connected component $\operatorname{Stab}(\overline{D}_{\ast})=G_{v_{\ast}}$ of $\overline{D}$. This concludes the proof. ∎